OpenStudy (anonymous):

no one is going to answer any more questions i understand because its time to go to sleep Zzz

OpenStudy (rogue):

Do you have a question?

OpenStudy (anonymous):

yes i have multiple would you help their geometry

OpenStudy (anonymous):

I can help.

OpenStudy (rogue):

I can try to help as well =)

OpenStudy (anonymous):

A reptile has a tail longer than its trunk. The length, in millimeters, of the tail and the trunk of the reptile was found to have a golden ratio. Which of these could be the tail and trunk lengths of the reptile? Answer Trunk – 55 mm, Tail – 89 mm Trunk – 55 mm, Tail – 144 mm Trunk – 89 mm, Tail – 90.16 mm Trunk – 89 mm, Tail – 160 mm

OpenStudy (anonymous):

i know its either the 3rd or 2nd

OpenStudy (rogue):

Its A.

OpenStudy (anonymous):

Golden ratio is about 1.6. Or you can be clever and remember how consecutive fibonacci numbers are in the golden ratio. 55 and 89 are two fibonacci numbers, so it's A.

OpenStudy (anonymous):

oh gosh really i googled it and people prior to me got the answer c or d

OpenStudy (rogue):

Shouldn't trust random strangers on the internet :P (but you can trust us! lol)

OpenStudy (anonymous):

you guys are awsome thanks willing to help with another .... :D

OpenStudy (anonymous):

Look at the triangle shown below. A triangle PQR is shown.DE is a segment joining point D on the side PQ and point E on the side QR.DE is parallel to PR.PQ has length 20 cm, PR has length 5 cm, DE has length 2 cm What is the length of the segment PD? Answer 8 cm 10 cm 12 cm 14 cm

OpenStudy (rogue):

Use similar triangles. Triangle DEG is similar to PRQ. \[\frac {QD}{DE} = \frac {QP}{PR}\]\[\frac {QD}{2} = \frac {20}{5}\]\[QD = 8 cm\]

OpenStudy (rogue):

PD = PQ - QD = 20 cm - 8 cm = 12 cm

OpenStudy (anonymous):

very nice thankyou :)

OpenStudy (rogue):

Your welcome :D

OpenStudy (anonymous):

The figure below shows a triangular lawn labeled PQR. There are two fences, AB and BC. What is the total length of the fence AB and BC? Answer 90 feet 130 feet 160 feet 200 feet